### v2

# Artin-Schreier towers of finite fields

Abstract.
Given a prime number $p$, we consider the tower of finite fields $\mathbb{F}_{p}=L_{-1}\subset L_{0}\subset L_{1}\subset\cdots$, where each step corresponds to an Artin-Schreier extension of degree $p$, so that for $i\geq 0$, $L_{i}=L_{i-1}[c_{i}]$, where $c_{i}$ is a root of $X^{p}-X-a_{i-1}$ and $a_{i-1}=(c_{-1}\cdots c_{i-1})^{p-1}$, with $c_{-1}=1$. We extend and strengthen to arbitrary primes prior work of Popovych for $p=2$ on the multiplicative order $O(c_{i})$ of the given generator $c_{i}$ for $L_{i}$ over $L_{i-1}$. In particular, for $i\geq 0$, we show that $O(c_{i})=O(a_{i})$, except only when $p=2$ and $i=1$, and that $O(c_{i})$ is equal to the product of the orders of $c_{j}$ modulo $L_{j-1}^{\times}$, where $0\leq j\leq i$ if $p$ is odd, and $i\geq 2$ and $1\leq j\leq i$ if $p=2$. We also show that for $i\geq 0$, the $\mathrm{Gal}(L_{i}/L_{i-1})$-conjugates of $a_{i}$ form a normal basis of $L_{i}$ over $L_{i-1}$. In addition, we obtain the minimal polynomial of $c_{1}$ over $\mathbb{F}_{p}$ in explicit form.
Key words and phrases:
Finite fields, Artin-Schreier extensions, multiplicative order
2010 Mathematics Subject Classification:
12E20,11T30
The first author was partially supported by grants from CONICET, FONCYT and SeCyT-UNCórdoba).
The second author was partially supported by NSERC grant 2023-00014
The third author was partially supported by NSERC grant 2020-04062
1. Introduction
We fix throughout the paper a prime number $p$ and an algebraic closure $E$ of the field $\mathbb{F}_{p}$ with $p$ elements. Recursively define $c_{-1},c_{0},c_{1},\dots\in E$ by taking $c_{-1}=1$ and selecting $c_{i+1}$ so that
$c_{i+1}^{p}-c_{i+1}-(c_{-1}\cdots c_{i})^{p-1}=0,\quad i\geq-1.$
We further set $L_{-1}=\mathbb{F}_{p}$, as well as
$L_{i+1}=L_{i}[c_{i+1}],\;a_{i}=(c_{-1}\cdots c_{i})^{p-1},\quad i\geq-1,$
and
$N_{i}=\frac{p^{p^{i+1}}-1}{p^{p^{i}}-1},\quad i\geq 0.$
Then
(1.1) $L_{-1}\subset L_{0}\subset L_{1}\subset L_{2}\subset\cdots,$
is a tower of finite fields, where $[L_{i+1}:L_{i}]=p$ for all $i\geq-1$ and
(1.2) $|L_{i}^{\times}|=p^{p^{i+1}}-1=(p-1)N_{0}\cdots N_{i},\quad i\geq-1.$
This construction was first given by Albert [A] in 1934, following prior work by Artin and Schreier [AS] in 1927. More general algebraic extensions of finite fields are discussed in the book [BS] by Brawley and Schnibben. When $p=2$, the above towers are described in Conway’s book [C] as well as in Lenstra’s notes [L].
The first extension $L_{0}/\mathbb{F}_{p}$ is connected to the period conjecture for the Bell numbers modulo $p$, which reduces to the question of whether or not the root $c_{0}$ of $X^{p}-X-1$ has multiplicative order $N_{0}$. This problem has been addressed, among others, by Wagstaff [Wa]; Car, Gallardo, Rahavandrainy, and Vaserstein [CGRV]; Montgomery, Nahm, and Wagstaff [MNW]; and Gallardo in [G1] and [G2]. While significant progress has been made, it remains open, as well as the question of whether $N_{0}$ is always square-free.
Albert’s construction was rediscovered by Ito, Kashiwara, and Huling [IKH]. They give an algorithm to produce irreducible polynomials of high degree related to the tower (1.1), and for $p=2$ they discuss the problem of finding primitive elements for each member of this tower. While finding the exact orders of given elements in a finite field may be too difficult, it is sometimes possible to find lower bounds for the orders of these elements. There is a vast literature on the subject and we refer the reader to the paper [BR] by Brochero Martínez and Reis, and the article [DMPS] by Dose, Mercuri, Pal, and Stirpe, as well as references therein.
One of the most recent papers studying the tower (1.1) in the case $p=2$ is due to Popovych [P]. He gives a lower bound for the order $O(c_{i})$ of $c_{i}$ as well as sufficient conditions for $O(c_{i})$, $i\geq 2$, to attain its largest possible value, which he proves to be $N_{1}\cdots N_{i}$. Under these conditions, he shows that one also has $O(a_{i})=O(c_{i})$, $i\geq 2$. As $O(c_{0})=3$ and the Fermat numbers $N_{i}$, $i\geq 1$, are relatively prime to $N_{0}=3$, it follows from (1.2) that for $i\geq 2$, $O(c_{i})=N_{1}\cdots N_{i}$ if and only if $c_{0}c_{i}$ is a primitive element of $L_{i}$. Questions of a similar nature had been addressed earlier by Wiedemann [W], who found other potentially primitive elements for $L_{i}$. However, his construction of the fields in the tower (1.1) was different, and an explicit relationship between the two constructions was left open. Seyfarth and Ranade [SR] show how to translate the problem of the primitivity of Wiedemann’s elements into a specific feature of the characteristic polynomial of certain matrices over $F_{2}$. Recursive constructions of elements of predictably high order in finite fields are useful for the development of discrete logarithm ciphers and pseudo-random number generators, and there are added benefits if the elements of high order determine normal bases over a ground field (see [PS]).
In this paper we extend Popovych’s results from [P] to all primes. In particular, Theorem 4.5 extends and strengthens [P, Theorem 4] by showing, unconditionally, that $O(c_{i})=O(a_{i})$ for all $i\geq 0$, except only when $p=2$ and $i=1$. In Corollary 4.7, which extends [P, Theorem 3], we give a formula for the common order of $c_{i}$ and $a_{i}$ as a product of the orders of $c_{j}$ modulo $L_{j-1}^{\times}$, where $0\leq j\leq i$ if $p$ is odd, and $i\geq 2$ and $1\leq j\leq i$ if $p=2$. Corollary 4.8 provides necessary and sufficient conditions for $O(c_{i})$ to attain its largest possible value, namely $N_{0}\cdots N_{i}$ if $p$ is odd, and $N_{1}\cdots N_{i}$ if $p=2$ and $i\geq 2$. As $p-1$ is relatively prime to all $N_{i}$ (see Lemma 3.1), it follows from (1.2) that if $p$ is odd, $i\geq 0$, and $z$ is a primitive element of $\mathbb{F}_{p}$, then $O(c_{i})=N_{0}\cdots N_{i}$ if and only if $zc_{i}$ is a primitive element of $L_{i}$. Corollary 4.9 extends [P, Corollary 2] and gives a lower bound for the order of $c_{i}$.
We begin the paper by reviewing Albert’s construction of Artin-Schreier field towers in Section 2 and establishing their basic properties in Section 3. Section 4 is devoted to the proof of Theorem 4.5 and its consequences. In Section 5 and Section 6 we establish two further properties of interest. Theorem 5.1 shows that for all $i\geq 0$, the $\mathrm{Gal}(L_{i}/L_{i-1})$-conjugates of $a_{i}$ form a normal basis of $L_{i}$ over $L_{i-1}$. Theorem 6.4 establishes a connection between the minimal polynomials of $c_{i}$ and $a_{i-1}$ over $\mathbb{F}_{p}$, $i\geq 1$, which is applied in Corollary 6.5 to explicitly exhibit the minimal polynomial of $c_{1}$ over $\mathbb{F}_{p}$.
2. Construction of the towers
Let us recall the following result due to Artin and Schreier.
Theorem 2.1.
Let $F$ be a field of prime characteristic $p$ and take $a\in F$. Then the polynomial $f=X^{p}-X-a\in F[X]$ is either irreducible or it has a root in $F$.
Proof.
Let $K$ be a splitting field for $f$ over $F$, and let $\alpha\in K$ be a root of $f$. As $F$ has characteristic $p$, we have
$(\alpha+1)^{p}-(\alpha+1)-1=(\alpha^{p}+1)-(\alpha+1)-1=\alpha^{p}-\alpha-1=0.$
It follows that the $p$ distinct elements $\alpha,\alpha+1,...,\alpha+(p-1)$ of $K$ are roots of $f$. Suppose $f$ has no roots in $F$, so that $\alpha\notin F$. Let $g\in F[X]$ be the minimal polynomial of $\alpha$ over $F$. The sum of the roots of $g$ in $K$ belongs to $F$. As $g|f$, this sum equals $d\alpha+k$, where $d$ is the degree of $g$ and $k\in F$. If $d<p$ then $\alpha\in F$. As this cannot be, it follows that $d=p$, so $g=f$ is irreducible. ∎
Lemma 2.2.
Let $F$ be a field of prime characteristic $p$ and take $a\in F$. Suppose $f=X^{p}-X-a$ has no roots in $F$ and set $K=F[\alpha]$, where $f(\alpha)=0$. Then the minimal and characteristic polynomial of $\alpha^{p-1}$ over $F$ is $X^{p}+(X-1)^{p-1}-a^{p-1}-1$. In particular, $\mathrm{tr}_{K/F}(\alpha^{p-1})=-1$.
Proof.
From $\alpha^{p}=\alpha+a$ we deduce $\alpha^{p-1}=1+\frac{a}{\alpha}.$ It follows that $F[\alpha]=F[\alpha^{p-1}]$, so $\alpha$ and $\alpha^{p-1}$ have the same degree over $F$, namely $p$ by Theorem 2.1. Moreover, $\alpha(\alpha^{p-1}-1)=a$ yields
$\alpha^{p-1}(\alpha^{p-1}-1)^{p-1}=a^{p-1},$
so the minimal polynomial and characteristic polynomial of $\alpha^{p-1}$ over $F$ is
$X(X-1)^{p-1}-a^{p-1}=(X-1+1)(X-1)^{p-1}-a^{p-1}=X^{p}+(X-1)^{p-1}-a^{p-1}-1.$
Note 2.3.
That $\mathrm{tr}_{K/F}(\alpha^{p-1})=-1$ can be seen directly by considering the map multiplication by $\alpha^{p-1}$, which has the following matrix of trace $-1$ relative to the basis $1,\alpha,...,\alpha^{p-1}$ of $K$ over $F$:
$\begin{bmatrix}0&a&0&\cdots&0&0\\ 0&1&a&\cdots&0&0\\ 0&0&1&\cdots&0&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&0&\cdots&1&a\\ 1&0&0&\cdots&0&1\\ \end{bmatrix}.$
Lemma 2.4.
We have
$\mathrm{tr}_{L_{i}/L_{-1}}(a_{i})=(-1)^{i+1},\quad i\geq-1.$
Proof.
By induction on $i$. The base case $i=-1$ is trivially true. Suppose that for some $i\geq-1$, we have
(2.1) $\mathrm{tr}_{L_{i}/L_{-1}}(a_{i})=(-1)^{i+1}.$
Since $\mathrm{tr}_{L_{i}/L_{-1}}(x^{p}-x)=0$ for all $x\in L_{i}$, Theorem 2.1 yields
$[L_{i+1}:L_{i}]=p.$
On the other hand, by Lemma 2.2, we have
(2.2) $\mathrm{tr}_{L_{i+1}/L_{i}}(c_{i+1}^{p-1})=-1.$
It follows from the transitivity of the trace, (2.1), and (2.2) that
$\mathrm{tr}_{L_{i+1}/L_{-1}}(a_{i+1})$ $=\mathrm{tr}_{L_{i}/L_{-1}}(\mathrm{tr}_{L_{i+1}/L_{i}}(a_{i+1}))$
$=\mathrm{tr}_{L_{i}/L_{-1}}(a_{i}\,\mathrm{tr}_{L_{i+1}/L_{i}}(c_{i+1}^{p-1}))$
$=\mathrm{tr}_{L_{i}/L_{-1}}(-a_{i})=(-1)^{i+2}.$
This proves Lemma 2.4. ∎
Theorem 2.1 and Lemma 2.4 imply that
(2.3) $[L_{i+1}:L_{i}]=p,\quad i\geq-1,$
which gives
(2.4) $|L_{i}|=p^{p^{i+1}},\quad i\geq-1.$
Lemma 2.5.
The following identity holds:
$c_{i}^{p^{p^{i}}}=c_{i}+(-1)^{i},\quad i\geq 0.$
Proof.
Consider the equations
$c_{i}^{p}-c_{i}=a_{i-1},$
$c_{i}^{p^{2}}-c_{i}^{p}=a_{i-1}^{p},$
$\vdots$
$c_{i}^{p^{p^{i}}}-c_{i}^{p^{p^{i}-1}}=a_{i-1}^{p^{p^{i}-1}}.$
Adding them up and appealing to Lemma 2.4, we obtain
$c_{i}^{p^{p^{i}}}-c_{i}=\mathrm{tr}_{L_{i-1}/L_{-1}}(a_{i-1})=(-1)^{i}.\qed$
It should be noted that the union, say $L$, of the members of the tower
$L_{-1}\subset L_{0}\subset L_{1}\subset\dots$
is the smallest subfield of $E$ that has no extensions of degree $p$. By construction, a basis for $L$ over $\mathbb{F}_{p}$ is formed by 1 and all elements of the form $c_{n_{1}}^{e_{1}}\cdots c_{n_{k}}^{e_{k}}$, where $k\geq 1$, $0\leq n_{1}<\dots<n_{k}$ and $1\leq e_{1},\dots,e_{k}<p$. We can also construct $L$ as the quotient of the polynomial ring $\mathbb{F}_{p}[X_{-1},X_{0},X_{1},\dots]$ by the ideal $I$ generated by
(2.5) $X_{-1}-1\text{ and }X_{i+1}^{p}-X_{i+1}-(X_{-1}\cdots X_{i})^{p-1},\quad i\geq-1.$
Indeed, there is a ring epimorphism $\mathbb{F}_{p}[X_{-1},X_{0},X_{1},\dots]\to L$ mapping $X_{j}\to c_{j}$ and containing (2.5) in its kernel. The resulting epimorphism from the quotient ring $\mathbb{F}_{p}[X_{-1},X_{0},X_{1},X_{2},\dots]/I$ to $L$ is seen to be injective by means of the preceding basis of $L$.
3. Basic properties
In this section we will establish some basic properties of the sequence of numbers $N_{i}$, $i\geq 0$, for a fixed prime $p$. These properties were noticed earlier by Popovych [P] in the case $p=2$; for odd primes the proofs are similar.
Recall that
$N_{i}=\frac{p^{p^{i+1}}-1}{p^{p^{i}}-1},\quad i\geq 0,$
so
(3.1) $\frac{p^{p^{i+1}}-1}{p-1}=N_{0}\cdots N_{i}.$
We can interpret $N_{i}$ in terms of the norm map $L_{i}\to L_{i-1}$, which, according to (2.4), is given by
$x\mapsto x^{1+p^{p^{i}}+p^{2p^{i}}+\cdots+p^{(p-1)p^{i}}}=x^{N_{i}}.$
Lemma 3.1.
The numbers $N_{i}$ and $p^{p^{i}}-1$ are relatively prime.
Proof.
Suppose, if possible, that $q$ is a prime factor of both $N_{i}$ and $p^{p^{i}}-1$. Then $p^{p^{i}}\equiv 1\mod q$ and therefore
$0\equiv N_{i}\equiv 1+p^{p^{i}}+\cdots+p^{p^{i}(p-1)}\equiv p\mod q,$
which implies $0\equiv 1\mod q$, a contradiction. ∎
Corollary 3.2.
Let $d$ be a positive divisor of $N_{i}$ different from 1. Then the order of $p$ modulo $d$ is precisely $p^{i+1}$. Thus, $p^{i+1}|\varphi(d)$. In particular, $p^{i+1}|\varphi(N_{i})$, and if $q$ is any prime divisor of $N_{i}$, then $q\equiv 1\mod p^{i+1}$, so that $d\equiv 1\mod p^{i+1}$.
Proof.
Since $d|N_{i}$, we have $p^{p^{i+1}}\equiv 1\mod d$. Therefore the order of $p$ modulo $d$ is a factor of $p^{i+1}$. This order cannot be a proper divisor of $p^{i+1}$, for otherwise $p^{p^{i}}\equiv 1\mod d$, which contradicts Lemma 3.1. ∎
Corollary 3.3.
If $i\neq j$, then $N_{i}$ and $N_{j}$ are relatively prime.
Proof.
Suppose $i>j$. By Lemma 3.1, $N_{i}$ is relatively prime to $p^{p^{i}}-1$ and hence to its factor $p^{p^{j+1}}-1$ and therefore to its divisor $N_{j}$. ∎
Lemma 3.4.
For $i\geq 0$, we have $c_{i}^{N_{i}}=a_{i-1}$.
Proof.
The minimal and characteristic polynomials of $c_{i}$ over $L_{i-1}$ are equal to $X^{p}-X-a_{i-1}$, so $a_{i-1}=N_{L_{i}/L_{i-1}}(c_{i})=c_{i}^{N_{i}}$. ∎
Lemma 3.5.
For $i\geq 0$, we have $c_{i}^{N_{0}\cdots N_{i}}=1=a_{i}^{N_{0}\cdots N_{i}}$.
Proof.
Since $c_{-1}\dots c_{i}\in L_{i}$, (2.4) gives $(c_{-1}\dots c_{i})^{p^{p^{i+1}}-1}=1.$ But $N_{0}\cdots N_{i}=\frac{p^{p^{i+1}}-1}{p-1}$ and $a_{i}=(c_{-1}\dots c_{i})^{p-1}$, so $a_{i}^{N_{0}\cdots N_{i}}=1$. If $i=0$, then Lemma 3.4 yields $c_{0}^{N_{0}}=a_{-1}=1$ as well as $c_{i+1}^{N_{0}\cdots N_{i}N_{i+1}}=a_{i}^{N_{0}\cdots N_{i}}=1.$
When $p=2$ a sharpening of Lemma 3.4 is possible. The next result follows from the fact that finite subgroups of the multiplicative group of a field are cyclic.
Lemma 3.6.
Let $K_{1}\subset K_{2}$ be an extension of fields. Suppose $a\in K_{2}^{\times}$ has finite order and set $m=\min\{n\geq 1\,|\,a^{n}\in K_{1}^{\times}\}$. Then $O(a)=m\times O(a^{m})$.
Lemma 3.7.
Suppose $p=2$. Then $c_{0}=a_{0}$ has order $N_{0}=3$, $O(c_{1})=N_{0}N_{1}=15$, $O(a_{1})=N_{1}=5$, and for $i\geq 2$, we have $c_{i}^{N_{1}\cdots N_{i}}=1=a_{i}^{N_{1}\cdots N_{i}}$.
Proof.
By Lemma 3.4, $c_{0}^{3}=1$. But $c_{0}^{2}=c_{0}+1$, so $c_{0}\neq 1$. Thus $O(c_{0})=O(a_{0})=3$. By (2.3), $[L_{1}:L_{0}]=2$. In particular, $c_{1}\notin L_{0}$. On the other hand, Lemma 3.4 gives $c_{1}^{5}=a_{0}=c_{0}\in L_{0}$. It follows from Lemma 3.6 that $O(c_{1})=15$. But $a_{1}=c_{1}c_{0}\notin L_{0}$ and $a_{1}^{5}=c_{1}^{5}c_{0}^{5}=c_{0}c_{0}^{2}=1$, so $O(a_{1})=5$.
We next show that $c_{i}^{N_{1}\cdots N_{i}}=1=a_{i}^{N_{1}\cdots N_{i}}$ for $i\geq 2$. Indeed, by Lemma 3.4, we have $c_{2}^{N_{1}N_{2}}=a_{1}^{N_{1}}=1$, so $a_{2}^{N_{1}N_{2}}=c_{2}^{N_{2}N_{1}}a_{1}^{N_{2}N_{1}}=1$. Suppose that for some $i\geq 2$, we have $c_{i}^{N_{1}\cdots N_{i}}=1=a_{i}^{N_{1}\cdots N_{i}}$. Then by Lemma 3.4,
$c_{i+1}^{N_{1}\cdots N_{i}N_{i+1}}=a_{i}^{N_{1}\cdots N_{i}}=1,$
which implies
$a_{i+1}^{N_{1}\cdots N_{i}N_{i+1}}=c_{i+1}^{N_{1}\cdots N_{i}N_{i+1}}a_{i}^{N_{1}\cdots N_{i}N_{i+1}}=1.\qed$
4. The orders of $c_{i}$ and $a_{i}$
Lemma 4.1.
For $i\geq 0$, we have $\gcd(N_{i}-(p^{p^{i}}-1),N_{0}\cdots N_{i})=1$.
Proof.
Suppose, if possible, that $q$ is a prime factor of both numbers. Then $q|N_{j}$ for some $0\leq j\leq i$. If $j<i$, then $q|(p^{p^{i}}-1)$. But $q|(N_{i}-(p^{p^{i}}-1))$, so $q|N_{i}$, against Corollary 3.3. If $j=i$, then from $q|(N_{i}-(p^{p^{i}}-1))$ we infer $q|(p^{p^{i}}-1)$, against Lemma 3.1. ∎
Lemma 4.2.
For $i\geq 0$ and $m\geq 1$, we have
$c_{i}^{m}\in L_{i-1}\Leftrightarrow c_{i}^{m(N_{i}-(p^{p^{i}}-1))}=a_{i-1}^{m}.$
Proof.
We have $|L_{i-1}|=p^{p^{i}}$ by (2.4) and $c_{i}^{N_{i}}=a_{i-1}$ by Lemma 3.4, so
$c_{i}^{m}\in L_{i-1}$ $\Leftrightarrow c_{i}^{mp^{p^{i}}}=c_{i}^{m}\Leftrightarrow c_{i}^{m(p^{p^{i}}-1)}=1$
$\Leftrightarrow c_{i}^{m(N_{i}-(N_{i}-(p^{p^{i}}-1)))}=1\Leftrightarrow a_{i-1}^{m}c_{i}^{-m(N_{i}-(p^{p^{i}}-1))}=1$
$\Leftrightarrow c_{i}^{m(N_{i}-(p^{p^{i}}-1))}=a_{i-1}^{m}.\qed$
Lemma 4.3.
For $i\geq 0$, let $m$ be the order of $c_{i}$ modulo $L_{i-1}^{\times}$. Then $m$ is relatively prime to the order of $a_{i-1}$.
Proof.
This is obvious if $i=0$, so we assume $i\geq 1$. We have $a_{i-1}^{N_{0}\cdots N_{i-1}}=1$ by Lemma 3.5, $\gcd(N_{0}\cdots N_{i-1},N_{i})=1$ by Corollary 3.3, and $m|N_{i}$ by Lemma 3.4. ∎
Proposition 4.4.
For $i\geq 0$, let $m$ be the order of $c_{i}$ modulo $L_{i-1}^{\times}$. Then
$O(c_{i})=m\times O(a_{i-1}).$
Proof.
By Lemma 3.5, we have $c_{i}^{N_{0}\cdots N_{i}}=1$, so Lemma 4.1 implies that $c_{i}$ and $c_{i}^{N_{i}-(p^{p^{i}}-1)}$ generate the same subgroup. In particular, $m$ is also the order of $c_{i}^{N_{i}-(p^{p^{i}}-1)}$ modulo $L_{i-1}^{\times}$. This, Lemma 3.6, Lemma 4.2, and Lemma 4.3 yield
$O(c_{i})=O(c_{i}^{N_{i}-(p^{p^{i}}-1)})=m\times O(a_{i-1}^{m})=m\times O(a_{i-1}).\qed$
Theorem 4.5.
If $i\geq 0$, then $O(c_{i})=O(a_{i})$, except only when $p=2$ and $i=1$.
Proof.
By definition, $a_{i}=c_{i}^{p-1}a_{i-1}$. Here $\gcd(p-1,N_{0}\cdots N_{i})=1$ by Lemma 3.1. If $i=0$, then Lemma 3.5 yields the desired result. Assume in what follows that $i\geq 1$. In this case, Lemma 3.5 yields that the orders of $c_{i}$ and $a_{i}$ modulo $L_{i-1}^{\times}$ are identical, say equal to $m$. Thus by Lemma 3.6,
(4.1) $O(a_{i})=m\times O(a_{i}^{m}).$
We have $a_{i}^{N_{0}\cdots N_{i}}=1$ by Lemma 3.5 and $\gcd(N_{i}-(p^{p^{i}}-1),N_{0}\cdots N_{i})=1$ by Lemma 4.1, so
(4.2) $O(a_{i}^{m})=O(a_{i}^{m(N_{i}-(p^{p^{i}}-1))}).$
Since $a_{i}^{m}=c_{i}^{(p-1)m}a_{i-1}^{m}$, Lemma 4.2 ensures that
(4.3) $a_{i}^{m(N_{i}-(p^{p^{i}}-1))}=c_{i}^{m(p-1)(N_{i}-(p^{p^{i}}-1))}a_{i-1}^{m(N_{i}-(p^{p^{i}}-1))}=a_{i-1}^{m((p-1)+N_{i}-(p^{p^{i}}-1))}.$
It follows from (4.2), (4.3), and Lemma 4.3 that
(4.4) $O(a_{i}^{m})=O(a_{i-1}^{(p-1)+N_{i}-(p^{p^{i}}-1)})=O(a_{i-1}^{(p-1)+N_{i}}),$
where the last equality follows from (3.1) and Lemma 3.5.
Suppose, if possible, that $q$ is a prime factor of $\frac{p^{p^{i}}-1}{p-1}$ and $(p-1)+N_{i}$. Then $p^{p^{i}}\equiv 1\mod q$, so $N_{i}\equiv p\mod q$. Thus $(p-1)+N_{i}\equiv 2p-1\mod q$ and therefore $2p\equiv 1\mod q$. This yields
$1\equiv(2p)^{p^{i}}\equiv 2^{p^{i}}p^{p^{i}}\equiv 2^{p^{i}}\mod q.$
Thus the order of 2 modulo $q$ is a positive power of $p$, whence $q\equiv 1\mod p$. In particular, $q>p$, so for any $\ell\geq 2$, we have $\ell q\geq 2q>2p>2p-1$. Since $2p\equiv 1\mod q$, this forces $2p-1=q$. But then
$1\equiv q\equiv 2p-1\equiv-1\mod p,$
whence $p=2$.
Suppose $p$ is odd. It follows from (3.1), Lemma 3.5, and the above analysis that
(4.5) $O(a_{i-1}^{(p-1)+N_{i}})=O(a_{i-1}).$
Combining Proposition 4.4 with (4.1), (4.4), and (4.5) we obtain the desired result.
Suppose next $p=2$. The cases $i=0$ and $i=1$ were discussed in Lemma 3.7. Suppose that $i\geq 2$. Then (4.4) becomes
(4.6) $O(a_{i}^{m})=O(a_{i-1}^{1+N_{i}-(2^{2^{i}}-1)})=O(a_{i-1}^{1+(2^{2^{i}}+1)-(2^{2^{i}}-1)})=O(a_{i-1}^{3}).$
Here $a_{i-1}^{N_{1}\cdots N_{i-1}}=1$ by Lemma 3.7, and $\gcd(N_{1}\cdots N_{i-1},3)=1$ by Lemma 3.3, so (4.6) implies
(4.7) $O(a_{i}^{m})=O(a_{i-1}).$
Combining Proposition 4.4 with (4.1) and (4.7) we obtain the sought result. ∎
Remark 4.6.
For $i\geq 0$, we write $M_{i}$ for the order of $c_{i}$ modulo $L_{i-1}^{\times}$. Since $[L_{i},L_{i-1}]=p$, it follows that $M_{i}>1$. By Lemma 3.4, we have $M_{i}|N_{i}$, whence $M_{i}\equiv 1\mod p^{i+1}$ by Corollary 3.2. If $p$ is odd, then $N_{i}\equiv 1\mod 2$, so in this case $M_{i}\equiv 1\mod 2p^{i+1}$ and therefore $M_{i}\geq 1+2p^{i+1}$. If $p=2$ and $i\geq 2$, then by the Euler-Lucas theorem, we have $M_{i}\equiv 1\mod 2^{i+2}$, which implies $M_{i}\geq 1+2^{i+2}$. If $p=2$, as the Fermat numbers $N_{i}$, $0\leq i\leq 4$, are prime, it follows that $M_{i}=N_{i}$ for all $0\leq i\leq 4$.
Corollary 4.7.
Let $i\geq 0$. If $p$ is odd, we have $O(c_{i})=M_{0}\cdots M_{i}$. If $p=2$, then $O(c_{0})=N_{0}$, $O(c_{1})=N_{0}N_{1}$, $O(c_{2})=N_{1}N_{2}$, $O(c_{3})=N_{1}N_{2}N_{3}$, $O(c_{4})=N_{1}N_{2}N_{3}N_{4}$, and for $i\geq 5$, $O(c_{i})=N_{1}N_{2}N_{3}N_{4}M_{5}\cdots M_{i}$.
Proof.
If $p$ is odd, the result follows immediately by repeatedly applying Proposition 4.4 and Theorem 4.5. If $p=2$, the result follows from Remark 4.6 and a repeated application of Proposition 4.4 and Theorem 4.5. ∎
Corollary 4.8.
Let $i\geq 0$. If $p$ is odd, then $O(c_{i})=N_{0}\cdots N_{i}$ if and only if $M_{j}=N_{j}$ for all $0\leq j\leq i$. If $p=2$ and $i\geq 5$, then $O(c_{i})=N_{1}\cdots N_{i}$ if and only if $M_{j}=N_{j}$ for all $5\leq j\leq i$.
Corollary 4.9.
Let $i\geq 0$. If $p$ is odd, then $O(c_{i})\geq(1+2p)\cdots(1+2p^{i+1})$. If $p=2$ and $i\geq 5$, then $O(c_{i})\geq N_{1}N_{2}N_{3}N_{4}(1+2^{7})\cdots(1+2^{i+2})$.
It is shown in [P, Theorem 5] by means of computer calculations that if $p=2$ then $M_{i}=N_{i}$ for all $5\leq i\leq 11$, whence $O(c_{i})=N_{1}\cdots N_{i}$ for all $2\leq i\leq 11$.
5. A normal basis for $L_{i}$ over $L_{i-1}$
Here we show that the Galois conjugates of $a_{i}$ form a normal basis for $L_{i}$ over $L_{i-1}$ for all $i\geq 1$. Given a field $K$ and $\alpha_{1},\dots,\alpha_{n}\in K$, we write $V(\alpha_{1},\dots,\alpha_{n})$ for the Vandormonde matrix in $M_{n}(K)$ associated to $(\alpha_{1},\dots,\alpha_{n})$, that is, with rows equal to $(1,\dots,1),(\alpha_{1},\dots,\alpha_{n}),\dots,(\alpha_{1}^{n-1},\dots,\alpha_{n}^{n-1})$.
Theorem 5.1.
For $i\geq 0$, let $\sigma_{i}$ be the generator of the Galois group of $L_{i}$ over $L_{i-1}$ given by $x\mapsto x^{p^{p^{i}}}$. Then $\{a_{i},{\sigma_{i}}(a_{i}),\dots,{\sigma_{i}^{p-1}}(a_{i})\}$ is a basis for $L_{i}$ over $L_{i-1}$.
Proof.
We know that $\{c_{i}^{p-1},\dots,c_{i},1\}$ is a basis of $L_{i}$ over $L_{i-1}$. Let $M$ be the matrix whose columns are the coordinates of $a_{i},{\sigma_{i}}(a_{i}),\dots,{\sigma_{i}^{p-1}}(a_{i})$ relative to this basis. We claim that $M$ is invertible. Indeed, by Lemma 2.5, we have
${\sigma_{i}}(c_{i})=c_{i}+(-1)^{i},{\sigma_{i}^{2}}(c_{i})=c_{i}+2(-1)^{i},\dots,{\sigma_{i}^{p-1}}(c_{i})=c_{i}+(p-1)(-1)^{i},$
and therefore
$a_{i}=a_{i-1}c_{i}^{p-1},{\sigma_{i}}(a_{i})=a_{i-1}(c_{i}+(-1)^{i})^{p-1},{\sigma_{i}^{2}}(a_{i})=a_{i-1}(c_{i}+2(-1)^{i})^{p-1},\dots,$
${\sigma_{i}^{p-2}}(a_{i})=a_{i-1}(c_{i}+(p-2)(-1)^{i})^{p-1},{\sigma_{i}^{p-1}}(a_{i})=a_{i-1}(c_{i}+(p-1)(-1)^{i})^{p-1}.$
Thus $M=DVS$, where $D=\mathrm{diag}(\binom{p-1}{0},\binom{p-1}{1},\dots,\binom{p-1}{p-1})$, $V$ is the Vandermonde matrix $V(0\times(-1)^{i},1\times(-1)^{i},\dots,(p-1)\times(-1)^{i})$, and $S=\mathrm{diag}(a_{i-1},\dots,a_{i-1})$.
6. Minimal polynomials
Theorem 6.1.
Let $K/F$ be a finite Galois extension with Galois group $G$. If $f\in F[X]$ is monic and irreducible and $g\in K[X]$ is a monic irreducible factor of $f$ in $K[X]$, then
$f=\underset{h\in Gg}{\Pi}h,$
where $Gg$ is the orbit of $g$ under the action of $G$. In particular, we have that $\deg f=\deg h\times|G|/|G_{g}|$, where $G_{g}$ is the stabilizer of $g$ in $G$.
Proof.
It is clear that $\underset{h\in Gg}{\Pi}h\in F[X]$ is a non-constant monic factor of $f$ in $K[X]$ and hence in $F[X]$. As $f$ is irreducible, $f=\underset{h\in Gg}{\Pi}h$. ∎
Theorem 6.2.
Let $K/F$ be a finite Galois extension with Galois group $G$. Then for $\alpha\in K$, we have
$f_{\alpha,F}=\underset{\beta\in G\alpha}{\Pi}(X-\beta).$
In particular, $\deg f=|G|/|G_{\alpha}|$ is a factor of $|G|$. Also in particular, if $F=F_{q}$ and $K=F_{q^{n}}$, then $G=\langle T\rangle$, where $T(x)=x^{q}$ for all $x\in K$, $T$ has order $n$, and if $m$ is the smallest positive integer such that $T^{m}(\alpha)=\alpha$, then $m|n$ and
$f_{\alpha,F}=(X-\alpha)\cdots(X-T^{m-1}(\alpha)).$
Proof.
Immediate consequence of Theorem 6.1. ∎
Theorem 6.3.
If $i\geq 0$, then $c_{i}$ and $a_{i}$ have degree $p^{i+1}$ over $\mathbb{F}_{p}$.
Proof.
By Theorem 6.2 the degree of $c_{i}$ over $\mathbb{F}_{p}$ is a factor of $p^{i+1}$. If the degree is not $p^{i+1}$, then necessarily $c_{i}^{p^{p^{i}}}=c_{i}$, which contradicts Lemma 2.5. Likewise, if the degree of $a_{i}$ is not $p^{i+1}$, then Lemma 2.5 yields
$c_{i}^{p-1}a_{i-1}=a_{i}=a_{i}^{p^{p^{i}}}=(c_{i}+(-1)^{i})^{p-1}a_{i-1},$
which implies $c_{i}^{p-1}=(c_{i}+(-1)^{i})^{p-1}$. It follows that $(c_{i}+(-1)^{i})c_{i}^{-1}\in\mathbb{F}_{p}$. But $(c_{i}+(-1)^{i})c_{i}^{-1}=1+(-1)^{i}c_{i}^{-1}$, which is not in $\mathbb{F}_{p}$ by (2.3). ∎
Theorem 6.4.
For $i\geq 1$, we have
$f_{c_{i},\mathbb{F}_{p}}(X)=f_{a_{i-1},\mathbb{F}_{p}}(X^{p}-X).$
In particular, this polynomial is irreducible in $\mathbb{F}_{p}[X]$.
Proof.
By definition, $f_{c_{i},L_{i-1}}(X)=X^{p}-X-a_{i-1}$. By Theorem 6.3, $a_{i-1}$ has exactly $p^{p^{i}}$ conjugates over $\mathbb{F}_{p}$, so Theorem 6.1 yields
$f_{c_{i},\mathbb{F}_{p}}(X)=(X^{p}-X-a_{i-1})(X^{p}-X-a_{i-1}^{p})\cdots(X^{p}-X-a_{i-1}^{p^{p^{i}-1}})=f_{{a_{i-1},\mathbb{F}_{p}}}(X^{p}-X).\qed$
Corollary 6.5.
The minimal polynomial of $c_{1}$ over $\mathbb{F}_{p}$ is equal to
$X^{p^{2}}+(X^{p}-X-1)^{p-1}-X^{p}-2.$
In particular, this polynomial is irreducible in $\mathbb{F}_{p}[X]$.
Proof.
We have $f_{c_{0},\mathbb{F}_{p}}(X)=X^{p}-X-1$, so $f_{c_{0}^{p-1},\mathbb{F}_{p}}(X)=X^{p}+(X-1)^{p-1}-2$ by Lemma 2.2. Thus, by Theorem 6.4,
$f_{c_{1},\mathbb{F}_{p}}(X)$ $=f_{c_{0}^{p-1},\mathbb{F}_{p}}(X^{p}-X)=(X^{p}-X)^{p}+(X^{p}-X-1)^{p-1}-2$
$=X^{p^{2}}+(X^{p}-X-1)^{p-1}-X^{p}-2.\qed$